OpenStudy (anonymous):

Anita and Joelle bowl together and their combined total score for one game was 425 points. Anita’s score was 40 less than twice Joelle’s. What were their scores? Write a system of equations to model the problem if x represents Joelle’s score and y represents Anita’s score. A. x – y = 425 y = 2x – 40 B. x + y = 425 y = 2x + 40 C. x + y = 40 y = 2x – 425 D. x + y = 425 y = 2x – 40

pooja195 (pooja195):

Let x = Anita's score and y = Joelle's score then x + y = 425 (equation 1) x = 2y - 40 (equation 2) . Use the definition of x from equation 2 and substitute it into equation 1: x + y = 425 2y - 40 + y = 425 3y - 40 = 425 3y = 465 y = 155 (Joelle's score) . Anita's score: x + y = 425 x + 155 = 425 x = 270 (Anita's score)